# On Viscosity Solutions of Hamilton-Jacobi Equations in the Wasserstein space and the Vanishing Viscosity Limit

### v1

@@ 3.2. Hopf-Lax Formula, Dynamic Programming Principle, regularity and infinitesimal behavior of optimizers
For the sake of presentation, we derive the Hopf-Lax formula from the associated control problem. Consider the value function considered in [BSOT]
$U_{K}(t,\mu)=\inf_{\zeta_{\cdot}\in\mathcal{A}_{[t,T]}(\mu)}{\int_{t}^{T}\mathcal{L}(\zeta_{s})ds+\mathcal{G}((\pi_{1})_{\#}\zeta_{T})},$
@@
$\mathcal{A}_{[t,T]}(\mu):=\Big\{\zeta_{\cdot}\in C([t,T];\mathcal{P}_{2}(\mathbb{R}^{2d})):\partial_{s}\mu_{s}+\text{div}(\text{Pr}(\zeta_{s})\mu_{s})=0,\,\mu_{s}=(\pi_{1})_{\#}\zeta_{s},\,\mu_{t}=\mu\,\,\,\forall s\in[t,T]\Big\}.$
@@
$U_{M}(t,\mu):=\inf_{(\mu_{\cdot},\alpha_{\cdot})\in\mathcal{A}^{M}_{[t,T]}(\mu)}{\int_{t}^{T}\int_{\mathbb{R}^{d}}L(\alpha_{s}(x))d\mu_{s}ds+\mathcal{G}(\mu_{T})},$
[…]
$\mathcal{A}^{M}_{[t,T]}(\mu):=\Big\{(\mu_{\cdot},\alpha_{\cdot})\in C([t,T];\mathcal{P}_{2}(\mathbb{R}^{2d})\times L^{2}_{\mu_{\cdot}}(\mathbb{R}^{d};\mathbb{R}^{d})):\partial_{s}\mu_{s}+\text{div}(\alpha_{s}\mu_{s})=0,\,\,\,s\in[t,T];\,\mu_{t}=\mu,\Big\}.$
@@
On the other hand, by Jensen inequality $\mathcal{L}(\zeta)\geq\mathcal{L}((\mathrm{Id}\times\Pr(\zeta))),\,\forall\zeta\in\mathcal{P}_{2}(\mathbb{R}^{2d})$. Therefore, $\forall(t,\mu)\in[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})$
$U_{K}(t,\mu)\geq\inf_{\zeta_{\cdot}\in\mathcal{A}_{[t,T]}(\mu)}{\int_{t}^{T}\mathcal{L}(\mathrm{Id}\times\Pr(\zeta_{s}))ds+\mathcal{G}((\pi_{1})_{\#}\zeta_{T})}\geq U_{M}(t,\mu).$
@@
$\mathcal{V}(t,\mu):$ $=\inf_{\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{d})}\left\{\mathcal{G}((\pi_{2})_{\#}(\gamma))+(T-t)\mathcal{L}\left(\frac{1}{T-t}\cdot(\pi_{1},\pi_{2}-\pi_{1})_{\#}\gamma\right)\right\}$
@@
$\mathcal{V}(t,\mu)=\inf_{\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{d})}\left\{\mathcal{G}(\exp_{\mu}(\gamma))+(T-t)\mathcal{L}\left(\frac{1}{T-t}\cdot\gamma\right)\right\}.$
@@
(3.6) $\lim_{t\to T^{-}}(T-t)\mathcal{L}\left(\frac{1}{T-t}\cdot\gamma\right)=\begin{cases}0\quad&\text{if }\gamma=\mu\otimes\delta_{o}\\ +\infty\quad&\text{otherwise}\end{cases}.$
Therefore, $\mathcal{V}(T,\mu)=\mathcal{G}(\exp(\mu\otimes\delta_{o}))=\mathcal{G}(\mu)$.
@@
Now, let $\gamma^{1\,3}\in\Gamma(\mu_{1},\mu_{3})$ be the composition of plans defined in [[AGS], Remark 5.3.3] as the plan given by
@@
Minimizing the LHS of the previous equation gives the desired result. ∎
@@
The function $\mathcal{V}$ is Lipschitz continuous in $0,\infty)\times\mathcal{P}_{2}(\mathbb{R}^{d})$, provided that ${\mathrm{Lip}}(\mathcal{G},W_{2})<\infty$. More precisely, for all $0<t<T$
$\mathrm{Lip}(\mathcal{V}(t,\cdot);W_{2})\leq\mathrm{Lip}(\mathcal{G};W_{2})$
@@
The proof is an adaptation of the proof of [Evans, Lemma 2, Chapter 3] to the Wasserstein framework.
@@
and pick $\theta\in\Gamma_{0}(\gamma,\nu).$ We note that $\xi:=(\pi_{3},\pi_{2}+\pi_{3}-\pi_{1})_{\#}\theta\in\mathcal{P}_{2,\nu}(\mathbb{R}^{d})$, and $\mathcal{L}(\zeta_{T-t})=\mathcal{L}(\frac{1}{T-t}\cdot\gamma)$, where for $\zeta_{T-t}$ we adopted the same notation used in the proof of the DPP 3.4. Therefore,
@@
${\underbrace{\leq}_{\zeta=\xi}}\mathcal{G}((\pi_{2})_{\#}\zeta)-\mathcal{G}(\exp_{\mu}(\gamma))+\varepsilon\leq\mathrm{Lip}(\mathcal{G};W_{2})W_{2}((\pi_{2})_{\#}\zeta,\exp_{\mu}(\gamma))+\varepsilon$
$=\mathrm{Lip}(\mathcal{G};W_{2})(\int_{\mathbb{R}^{3d}}|z-x|^{2}d\theta(x,y,z))^{\frac{1}{2}}+\varepsilon$
@@
We now prove the time Lipschitz estimate. By the DPP, Proposition 3.4, fixing $0<t<t+h<T$ for $h>0$ and choosing $\gamma=(\mu\otimes\delta_{o})\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ we have
(3.12) $\mathcal{V}(t,\mu)\leq\mathcal{V}(t+h,\mu)+h\mathcal{L}\left(\frac{1}{h}\cdot(\mu\otimes\delta_{o})\right)=\mathcal{V}(t+h,\mu),$
@@
In particular, we have by the one sided Cauchy-Schartz inequality
@@
If $\text{Lip}(\mathcal{G};W_{2})<\infty$, then there exists a constant $C>0$, depending only on $\text{Lip}(\mathcal{G};W_{2})$ and on the coercitivity parameters of $L$, s.t.
@@
Moreover, by the coercitivity assumption on the Lagrangian,
@@
with $C$ depending on $\mathcal{G}$ and on the growth parameters of $L$, and using the subadditivity of the square root, we conclude the claim.
@@ 3.3. Existence result
We show now that the Hopf-Lax formula gives a viscosity solution in the sense of Definition 2.16 for equation (1.1).
@@ 3.3.1. Subsolution property
Fix $(t,\mu)\in(0,T)\times\mathcal{P}_{2}(\mathbb{R}^{d})$ and let $(r,\gamma)\in\partial^{+}U(t,\mu)$, namely $\forall\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$
$U(t+h,\exp_{\mu}(h\cdot\xi))-U(t,\mu)\leq h\big<\gamma,\xi\big>_{\theta,\mu}+rh+o(h\|\xi\|_{\mu})+o(h),\quad\forall h>0\text{ small enough},$
where $\theta\in\Gamma_{\mu}(\gamma,\xi).$ By the DPP, we have
$U(t,\mu)\leq U(t+h,\exp_{\mu}(\widetilde{\xi}))+h\mathcal{L}\left(\frac{1}{h}\cdot\widetilde{\xi}\right)\quad\forall\widetilde{\xi}\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}).$
@@
$0\leq h\mathcal{L}(\xi)+h\big<\gamma,\xi\big>_{\theta,\mu}+rh+o(h)+o(h\|\xi\|_{\mu}),\quad\forall\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}),h>0\text{ small enough}.$
@@
$0\leq\mathcal{L}(\xi)+\big<\gamma,\xi\big>_{\theta,\mu}+r.$
Therefore, since $\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ is arbitrary, we can invoke Lemma 3.1 to obtain
@@ 3.3.2. Supersolution Property
Fix $(t,\mu)\in(0,T)\times\mathcal{P}_{2}(\mathbb{R}^{d})$ and let $(r,\gamma)\in\partial^{-}U(t,\mu)$. Therefore, $\forall\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ we have
$U(t+h,\exp_{\mu}(\xi))-U(t,\mu)\geq\big<\gamma,\xi\big>_{\theta,\mu}+rh+o(\|\xi\|_{\mu})+o(h),\quad\forall h>0\text{ small enough.}$
@@
$U(t,\mu)+h\geq U(t+h,\exp_{\mu}(\xi))+h\mathcal{L}\left(\frac{1}{h}\cdot\xi\right).$
@@
We conclude by dividing by $h>0$ and passing to the limit as $h\to 0$.

### v2

@@ 3.2. Hopf-Lax Formula, Dynamic Programming Principle, regularity and infinitesimal behavior of optimizers
For the sake of presentation, we derive the Hopf-Lax formula from the associated control problem. Consider the value function introduced in [9]
$U_{K}(t,\mu):=\inf_{\zeta_{\cdot}\in\mathcal{A}_{[t,T]}(\mu)}{\int_{t}^{T}\mathcal{L}(\zeta_{s})ds+\mathcal{G}((\pi_{1})_{\#}\zeta_{T})},$
@@
$\mathcal{A}_{[t,T]}(\mu):=\Big\{\zeta_{\cdot}\in C([t,T];\mathcal{P}_{2}(\mathbb{R}^{2d})):\partial_{s}\mu_{s}+\text{div}(\text{Pr}(\zeta_{s})\mu_{s})=0,\,\mu_{s}=(\pi_{1})_{\#}\zeta_{s}\,\,\,\forall s\in[t,T],\,\mu_{t}=\mu\Big\}.$
@@
$U_{M}(t,\mu):=\inf_{(\mu_{\cdot},\alpha_{\cdot})\in\mathcal{A}^{M}_{[t,T]}(\mu)}{\int_{t}^{T}\int_{\mathbb{R}^{d}}L(\alpha_{s}(x))d\mu_{s}(x)ds+\mathcal{G}(\mu_{T})},$
where for $t<T$ and $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$,
$\mathcal{A}^{M}_{[t,T]}(\mu):=\Big\{(\mu_{\cdot},\alpha_{\cdot})\in C([t,T];\mathcal{P}_{2}(\mathbb{R}^{2d})\times L^{2}_{\mu_{\cdot}}(\mathbb{R}^{d};\mathbb{R}^{d})):\partial_{s}\mu_{s}+\text{div}(\alpha_{s}\mu_{s})=0\,\,\,\forall s\in[t,T],\,\mu_{t}=\mu\Big\}.$
@@
On the other hand, by Jensen inequality $\mathcal{L}(\zeta)\geq\mathcal{L}((\mathrm{Id}\times\Pr(\zeta))_{\#}((\pi_{1})_{\#}\zeta)),\,\forall\zeta\in\mathcal{P}_{2}(\mathbb{R}^{2d})$. Therefore, $\forall(t,\mu)\in[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})$
$U_{K}(t,\mu)\geq\inf_{\zeta_{\cdot}\in\mathcal{A}_{[t,T]}(\mu)}{\int_{t}^{T}\mathcal{L}((\mathrm{Id}\times\Pr(\zeta_{s}))_{\#}\mu_{s})ds+\mathcal{G}((\pi_{1})_{\#}\zeta_{T})}\geq U_{M}(t,\mu).$
@@
$\mathcal{V}(t,\mu):$ $=\inf_{\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})}\left\{\mathcal{G}((\pi_{2})_{\#}(\gamma))+(T-t)\mathcal{L}\left(\frac{1}{T-t}\cdot(\pi_{1},\pi_{2}-\pi_{1})_{\#}\gamma\right)\right\}$
@@
$\mathcal{V}(t,\mu)=\inf_{\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})}\left\{\mathcal{G}(\exp_{\mu}(\gamma))+(T-t)\mathcal{L}\left(\frac{1}{T-t}\cdot\gamma\right)\right\}.$
@@
$\lim_{t\to T^{-}}(T-t)\mathcal{L}\left(\frac{1}{T-t}\cdot\gamma\right)=\begin{cases}0\quad&\text{if }\gamma=\mu\otimes\delta_{0}\\ +\infty\quad&\text{otherwise}\end{cases}.$
Therefore, $\mathcal{V}(T,\mu)=\mathcal{G}(\exp_{\mu}(\mu\otimes\delta_{0}))=\mathcal{G}(\mu)$.
@@
Now, let $\gamma^{1\,3}\in\Gamma(\mu_{1},\mu_{3})$ be the composition of plans defined in [2, Remark 5.3.3] as the plan given by
@@
Since $\varepsilon$ is arbitrary, we have the claim.
@@
The function $\mathcal{V}$ defined in (3.2) is Lipschitz continuous in $0,\infty)\times\mathcal{P}_{2}(\mathbb{R}^{d})$, provided that ${\mathrm{Lip}}(\mathcal{G};W_{2})<\infty$. More precisely, for all $0<t<T$
$\mathrm{Lip}(\mathcal{V}(t,\cdot);W_{2})\leq\mathrm{Lip}(\mathcal{G};W_{2}),$
@@
The proof is an adaptation of the proof of [32, Lemma 2, Chapter 3] to the Wasserstein framework.
@@
and pick $\theta\in\Gamma_{0}(\gamma,\nu).$ We note that $\xi:=(\pi_{3},\pi_{2}+\pi_{3}-\pi_{1})_{\#}\theta\in\mathcal{P}_{2,\nu}(\mathbb{R}^{d})$, and $\mathcal{L}(\xi_{T-t})=\mathcal{L}\left(\frac{1}{T-t}\cdot\gamma\right)$, where for $\xi_{T-t}$ we adopted the same notation used in the proof of the DPP, Proposition 3.4. Therefore,
@@
${\underbrace{\leq}_{\zeta=\xi}}\mathcal{G}((\pi_{2})_{\#}\xi)-\mathcal{G}(\exp_{\mu}(\gamma))+\varepsilon\leq\mathrm{Lip}(\mathcal{G};W_{2})W_{2}((\pi_{2})_{\#}\xi,\exp_{\mu}(\gamma))+\varepsilon$
$=\mathrm{Lip}(\mathcal{G};W_{2})\left(\int_{\mathbb{R}^{3d}}|z-x|^{2}d\theta(x,y,z)\right)^{\frac{1}{2}}+\varepsilon$
@@
We now prove the time Lipschitz estimate. By the DPP, Proposition 3.4, fixing $0<t<t+h<T$ for $h>0$ and choosing $\gamma=(\mu\otimes\delta_{0})\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ we have
(3.15) $\mathcal{V}(t,\mu)\leq\mathcal{V}(t+h,\mu)+h\mathcal{L}\left(\frac{1}{h}\cdot(\mu\otimes\delta_{0})\right)=\mathcal{V}(t+h,\mu),$
@@
In particular, we have by the one sided Cauchy Schwarz inequality
@@
If $\mathrm{Lip}(\mathcal{G};W_{2})<\infty$, then there exists a constant $C>0$, depending only on $\mathrm{Lip}(\mathcal{G};W_{2})$ and on the coercivity parameters of $L$, s.t.
@@
Moreover, by the coercivity assumption on the Lagrangian,
@@
with $C$ depending on $\mathcal{G}$ and on the growth parameters of $L$. Using the subadditivity of the square root, we conclude the claim.
@@ 3.3. Existence result
We now show that the Hopf–Lax formula defines a viscosity solution, in the sense of Definition 2.19, to equation (1.1).
@@ 3.3.1. Subsolution property
Fix $(t,\mu)\in(0,T)\times\mathcal{P}_{2}(\mathbb{R}^{d})$ and let $(r,\gamma)\in\partial^{+}\mathcal{V}(t,\mu)$, namely $\forall\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ and $\theta\in\Gamma_{\mu}(\gamma,\xi)$
$\mathcal{V}(t+h,\exp_{\mu}(h\cdot\xi))-\mathcal{V}(t,\mu)\leq h\big<\gamma,\xi\big>_{\theta,\mu}+rh+o(h\|\xi\|_{\mu})+o(h),\quad\forall h>0\text{ small enough}.$
By the DPP, Proposition 3.4, we have
$\mathcal{V}(t,\mu)\leq\mathcal{V}(t+h,\exp_{\mu}(\widetilde{\xi}))+h\mathcal{L}\left(\frac{1}{h}\cdot\widetilde{\xi}\right)\quad\forall\widetilde{\xi}\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}).$
@@
$0\leq h\mathcal{L}(\xi)+h\big<\gamma,\xi\big>_{\theta,\mu}+rh+o(h)+o(h\|\xi\|_{\mu}),\quad\forall\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}),\theta\in\Gamma_{\mu}(\gamma,\xi),h>0\text{ small enough}.$
@@
(3.18) $0\leq\mathcal{L}(\xi)+\big<\gamma,\xi\big>_{\theta,\mu}+r.$
Therefore, since $\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ and $\theta\in\Gamma_{\mu}(\gamma,\xi)$ are arbitrary, we can invoke Lemma 3.1 to obtain
@@
Remark 3.7.
The proof of the subsolution property motivates our choice of superdifferential. Indeed, in general, the Fenchel duality, stated in equation (3.1) of Lemma 3.1, does not appear to hold when the supremum is restricted to the tangent space. More precisely, for a fixed $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, and for any $\gamma\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$, we only have
(3.19) $-\mathcal{H}(\gamma)$ $\leq-\sup_{\xi\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})}\Big\{-\langle\xi,\gamma\rangle_{\mu}-\mathcal{L}(\xi)\Big\}$
$=\inf_{\xi\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})}\Big\{\langle\xi,\gamma\rangle_{\mu}+\mathcal{L}(\xi)\Big\}.$
At present, we do not know whether equality holds for a general Lagrangian under our standing assumptions. Nevertheless, under the additional invariance property for the Lagrangian
(3.20) $\mathcal{L}(\mathrm{P}(\xi))\leq\mathcal{L}(\xi),$
the inequality in (3.19) is indeed an equality as proven in Lemma 3.1. In this case, we could have used, in the definition of viscosity subsolution, the Kantorovich superdifferential, see Definition 2.4, intersected with the tangent space $\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$. Indeed, the inequality (3.18) would read as
$0\leq\mathcal{L}(\xi)+\big<\gamma,\xi\big>_{\mu}+r,\quad\forall\xi\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}),$
and, under the decreasing property (3.20) and by Lemma 3.1,
$0\leq-\mathcal{H}(\gamma)+r,$
concluding the argument.
We underline that this condition is reminiscent of the decreasing property along a projection operator introduced in [36, Equation (5.7), p. 27], where the projection is $\text{Pr}_{\mu}\circ\text{P}:\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})\to\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$, see [36, Definition  2.5].
@@ 3.3.2. Supersolution Property
Fix $(t,\mu)\in(0,T)\times\mathcal{P}_{2}(\mathbb{R}^{d})$ and let $(r,\gamma)\in\partial^{-}\mathcal{V}(t,\mu)$. Therefore, $\forall\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ we have
$\mathcal{V}(t+h,\exp_{\mu}(\xi))-\mathcal{V}(t,\mu)\geq\big<\gamma,\xi\big>_{\theta,\mu}+rh+o(\|\xi\|_{\mu})+o(h),\quad\forall h>0\text{ small enough.}$
@@
$\mathcal{V}(t,\mu)+h\geq\mathcal{V}(t+h,\exp_{\mu}(\xi))+h\mathcal{L}\left(\frac{1}{h}\cdot\xi\right).$
@@
We conclude dividing by $h>0$ and passing to the limit as $h\to 0$.
Remark 3.8.
We leave for future work the analysis of the geodesic semiconcavity regularization property mentioned at the beginning of this section, in the case where the Lagrangian satisfies, for some $\theta>0$, the additional estimate
$\frac{1}{2}L(v_{1})+\frac{1}{2}L(v_{2})\leq L\!\left(\frac{v_{1}+v_{2}}{2}\right)+\frac{1}{\theta}\lvert v_{1}-v_{2}\rvert^{2},\qquad\forall\,v_{1},v_{2}\in\mathbb{R}^{d},$
as in [32, Lemma 4, p. 131]. Another interesting direction for future research is to establish geodesic semiconcavity in the Wasserstein space via PDE-based estimates for viscosity solutions of both equations (1.1) and (1.2).